let t be a real number such that t^2=at+b for some positive integers a and b. then for any choice of positive integers a and b, t^3 is never equal to?

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
This question can be answered only if options are given.
As thre can be infinite answers.
Please repost the question with options or you can also attach an image
Vikas TU
4 years ago
Dear student 
t^2 = at + b 
t^3 = at^2 + bt 
t^3 = a(at+b)+ bt
t^3 = (a^2+b)t + ab 
Check the option now. 
Hope this will help